36. What is the price of a stock expected to be in 4 years if the stock is expected to pay a dividend every year forever, the expected return for the stock is 15.61 percent per year, the next dividend is expected in 1 year, the next dividend is expected to be 8.21 dollars, and all subsequent dividends are expected to grow by 4.19 percent per year?
Answer #1

The expected price of the stock in 4 years = Dividend in year 5/(r - g)

Dividend in year 5 = D1 * (1 + g)^4

Dividend in year 5 = 8.21 * (1 + 0.0419)^4 = 9.6749183655

The expected price of the stock in 4 years = 9.6749183655/(0.1561 - 0.0419)

The expected price of the stock in 4 years = $84.7190750044

Can you please upvote? Thank You :-)